Given the following production function; Y=AX^(3/4) X^(1/4)

Required:
Calculate the degree of homogeneity of the above function and comment
on the returns to scale.
1
Expert's answer
2019-10-21T12:33:46-0400

Solution:

We are going to Calculate the degree of homogeneity of a function Y = "A \space X^{(\frac {3} {4})} X^{(\frac {1} {4})}"


We know, If the bases are equal in the multiplication, we can add the powers


So, the given equation can be written as "Y =A \space X^{(\frac {3} {4})+ (\frac {1} {4})}" = "A \space X^{(\frac {4} {4})} = A X"


Let f (X) = Y = A X


f(kX) = A (k X) = "k^{1}" (A X) = "k^{1} f(X)"

Yes, it is a homogeneous function of degree = 1

Measure of returns to scale is 1


Answer: The given function is homogeneous function of order 1